OpenStudy (anonymous):

Finn bought a yoyo from a company that claims that, with each retraction, the string rolls up by 80% of the original length. He sets up a tape measure and throws the yoyo 3 times. His data are charted below. Throw Length of string (feet) 1 3 2 2.4 3 1.92 Finn wants to find the sum of the length of string after 10 throws. What is the sum of the lengths, rounded to the nearest hundredth?

OpenStudy (anonymous):

13.39 feet 15.00 feet <--my answer 4.46 feet 0.40 feet

OpenStudy (mathmate):

With initial length of \(a_1\)=3 feet, and common ratio r= 0.8, the total length of n throws is given by the sum: \(S(n)=a_1(1-r^n)/(1-r)\) For example, after 1 throw, n=1, S(1)=3(1-0.8)/(1-0.8)=3 feet After 2 throws, \(S(2)=3(1-0.8^2)/(1-0.8)=5.4 feet.\) You can then calculate S(10) but putting n=10.

OpenStudy (anonymous):

So, wasn't I right then? :)

zepdrix (zepdrix):

The example mathmate gave,\[\large\rm S_2=3\cdot\frac{1-(0.8)^2}{1-0}\]As was mentioned though, we want the sum of the first 10,\[\large\rm S_{10}=3\cdot\frac{1-(0.8)^{10}}{1-0.8}\]So ummmm, yes maybe you are right :o But our your calculator.

zepdrix (zepdrix):

blah typos... bust out your calculator*

OpenStudy (anonymous):

2.092 is what I got :)

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hmm you shouldn't get less than 3 -_-

zepdrix (zepdrix):

\[\large\rm S_{10}=3*((1-.8\wedge(10))\div(1-.8))\]

OpenStudy (anonymous):

ok :) Hold on.

OpenStudy (anonymous):

That's so weird! Now I got 2.077?? My calculator must be broken..

OpenStudy (anonymous):

Ok, so for the first part I got 3.32212255

OpenStudy (anonymous):

Then 1.25 for the second part

zepdrix (zepdrix):

uhhhhh 0_o

zepdrix (zepdrix):

(1-.8^(10))/(1-.8) = 4.463129... And then multiply by 3. I dunno what you're doing there :d weird

OpenStudy (anonymous):

13.389 :)

OpenStudy (anonymous):

Thanks<3!!!